# Variational Mode Decomposition and Linear Embeddings are What You Need For Time-Series Forecasting

### v1

@@ 1 Introduction
Be that as it may, accomplishing optimal results with machine learning models includes a few preprocessing steps such as data cleaning, data decomposition, hyperparameter tuning, and etc. Choosing the proper time-series expectation calculation from a plenty of existing strategies is additionally significant.
One broadly related algorithm in time series forecasting is the Transformer, that is well known for its performance in Natural Language Processing (NLP), speech recognition, and computer vision assignments. However, Ailing Zeng et al has raised questions about the adequacy of Transformers in time-series forecast, inciting comparative execution assessments against easier linear-based calculations such as Linear, NLinear, and DLinear. Comparative tests have appeared that direct models frequently outperformed Transformer-based models like FEDFormer, AutoFormer, Informer, LogTrans, and PyraFormer in terms of Mean Squared Error (MSE) and Mean Absolute Error (MAE) (Zeng et al., 2022).
In expansion to algorithm choice, data decomposition strategies essentially affect forecast results. Variational Mode Decomposition (VMD) is one such strategy competent of lessening data instability (Dragomiretskiy & Zosso, 2014). The performance of VMD in time-series information forecast was highlighted in a 2021 study about comparing Recurrent Neural Network (RNN)-based models under different scenarios with PM2.5 air quality data over a few cities in China: 1) without decomposition method, 2) using Empirical Mode Decomposition (EMD), and 3) with VMD utilization. Based on the results, it was shown that VMD consistently achieved the least Root Mean Squared Error (RMSE), especially with BLSTM model (Zhang et al., 2021).
Given VMD’s capability to diminish data volatility and good performance of linear-based models, there exists potential for more compelling time-series data forecasting indeed with less stable data. These models will be compared against other deep learning models such as RNN, Long Short Term Memory (LSTM), and Bidirectional LSTM (BLSTM), which are less difficult however broadly utilized in time-series forecasting researches.
Propelled by the given reasons, this think about points to examine ”Variational Mode Deterioration with Straight Embeddings for Successful Time-Series Information Expectation.” By utilizing VMD for data decomposition and assessing embarrassingly simple linear models (Linear, NLinear, and DLinear) next to RNN-based models, the is to develop a predictive model competent of accomplishing minimal error rates indeed with unstable data conditions.
@@ 2.1 Time-Series Forecasting Models
Large Language Models. Large Language Models are models that were designed to understand, work, and generate with various human language. However, it was also recently developed to work with images for Computer Vision (CV) by utilizing its robustness in pattern recognition and reasoning. However, well-known LLM such as GPT and BERT can be used in forecasting time series data by fine-tuning it based on temporal data and also unifying their framework so that it can be used in diverse sequential task (classification, anomaly detection, forecasting, etc.) (Zhou et al., 2023). Time-LLM were developed in this year, a time series LLM model were developed to work with time-series data by reprogramming time series inputs into text prototype representations and guiding LLM’s reasoning by augmenting the input context with declarative prompts (Jin et al., 2024). It was proven competitive against many recently advanced time series forecasting method such as GPT4TS, DLinear, PatchTST, etc.
@@ 2.2 Data Mode Decomposition
Data mode decomposition is a technique used to decompose complex time series data into many data modes or Intrinsic Mode Function (IMF) with less complexity. Empirical Mode Decomposition (EMD) is one of data decomposition method that decomposes a signal into a set of intrinsic mode functions (IMFs) and a residual trend, capturing components of varying frequency ranges adaptively. EMD operates by iteratively sifting the data to identify and extract these IMFs, making it particularly effective for non-stationary and nonlinear time series data (Huang et al., 1998). However, EMD can be sensitive to noise and may suffer from mode mixing, where different modes are incorrectly combined. To address those limitations, Variational Mode Decomposition (VMD) was introduced as a more robust data decomposition method. VMD decomposes a signal into modes by solving a variational problem that seeks to minimize the bandwidth of each mode while ensuring they are as distinct as possible from each other. It utilized Lagrange function and Alternate Direction Method of Multipliers (ADMM) (Dragomiretskiy & Zosso, 2014). Lagrange function uses $\lambda$ penalty term to enforce constrain for decomposed components. Lagrange is a quite complex function that can be simplified by implementing ADMM to split the function into several vital parts of function. ADMM works by iterating those functions until it reaches convergence (Boyd et al., 2011). This method is less prone to noise and mode mixing issues compared to EMD, as it uses a principled optimization approach to separate modes in the frequency domain. Dynamic Mode Decomposition (DMD) is an another data decomposition method that is capable of extracting important dynamic characteristics of the data. DMD decomposes the data into several data modes or IMFs by utilizing matrix analysis and singular value decomposition (Proctor et al., 2016).
@@ 3.1 VMD-Linear Framework
The structure of VMD-Linear shown in Figure 1 was inspired by VMD-BLSTM framework introduced by Zhang et al. (2021). The goal is to reduce data bandwidth in all IMF or data modes with the sum of all IMF equals to that of the original data. LTSF-Linear models will be assigned on each IMF for time-series forecasting. For the final step, all forecast results will be summed together to get the final forecast result.
@@ 3.2 Variational Mode Decomposition
The symbol $\phi_{k}(t)$ is a non-decreasing function (can be considered a wave) with the condition that $\phi_{k}(t)\geq 0$, and $A_{k}(t)$ is the envelope of the wave $\phi_{k}(t)$. All IMF formed must satisfy Equation 2.1. VMD applies Hilbert transform to obtain the analytic frequency spectrum ($Au_{k}$) of each mode $u_{k}$. In the transformation process, each mode is multiplied by the center frequency ($e^{-jw_{k}t}$) to shift its frequency spectrum towards the ”baseband” (Dragomiretskiy & Zosso, 2014).
@@ 3.3 Long-term Time Series Forecasting Linear
NLinear NLinear model is used to anticipate distribution shifts in the data by reducing the last input passed to the Linear layer and then adding back the reduced input before making the final prediction.
@@
DLinear DLinear enhances the simple linear model by utilizing moving average to split the data into Trend and Seasonal with each going through their seperate linear layer. This model should be used if there is a clear trend in the data.
@@ 4.1 Dataset
    Wind Turbine, The following data contains information about a certain wind turbine in Turkey that were captured using the SCADA system. Wind Turbine related data such as wind speed, wind direction, and power generated were captured in 10-minute intervals with a total of 50,500 data points.
@@
    SEA Air Quality Dataset We also used various real-world Air Quality Index (AQI) datasets from many cities recorded from 27 January 2022 to 22 Mei 2023. Those cities were Central Singapore, Banjarbaru, Batam, Jakarta, South Jakarta, Jambi, Malang, Medan, Samarinda, and Semarang.
@@ 4.2 Experimental Protocols
Dataset Lin DLin NLin LSTM BiLSTM RNN
@@
Dataset Lin DLin NLin LSTM BiLSTM RNN

### v2

@@ 1 Introduction
Achieving optimal results with machine learning models requires a several preprocessing steps such as data cleaning, data decomposition, hyperparameter tuning, and etc. Additionally, selecting the appropriate time-series forecasting method from a the available strategies is crucial.
One widely used algorithm in time series forecasting is the Transformer, that is well known for its performance in Natural Language Processing (NLP), speech recognition, and computer vision assignments. However, Zeng et al. (2022) has raised questions about the adequacy of Transformers in time-series forecast, inciting comparative execution assessments against easier linear-based calculations such as Linear, NLinear, and DLinear. Comparative tests have appeared that direct models frequently outperformed Transformer-based models like FEDFormer, AutoFormer, Informer, LogTrans, and PyraFormer in terms of Mean Squared Error (MSE) and Mean Absolute Error (MAE) (Zeng et al., 2022).
Furthermore, data decomposition strategies essentially affect forecast results. Variational Mode Decomposition (VMD) is one such strategy competent of lessening data instability (Dragomiretskiy & Zosso, 2014). The performance of VMD in time-series information forecast was highlighted in a 2021 study about comparing Recurrent Neural Network (RNN)-based models under different scenarios with PM2.5 air quality data over a few cities in China: 1) without decomposition method, 2) using Empirical Mode Decomposition (EMD), and 3) with VMD utilization. Based on the results, it was shown that VMD consistently achieved the least Root Mean Squared Error (RMSE), especially with Bidirectional LSTM model (Zhang et al., 2021).
Given VMD’s capability to diminish data volatility and good performance of linear-based models, there exists potential for more compelling time-series data forecasting indeed with less stable data. These models will be compared against other deep learning models such as RNN, Long Short Term Memory (LSTM), and Bidirectional LSTM (BiLSTM), which are less difficult however broadly utilized in time-series forecasting researches.
motivated by the reasons outlined, this aims to investigate ”Variational Mode Deterioration with Straight Embeddings for Successful Time-Series Information Expectation.” The goal of employing VMD for data decomposition and evaluating simplistic linear models (Linear, NLinear, and DLinear) alongside RNN-based models is to develop a predictive framework capable of achieving minimal error rates even under volatile data conditions.
@@ 2.1 Time-Series Forecasting Models
Large Language Models. Large Language Models are models that were designed to understand, work, and generate with various human language. However, it was also recently developed to work with images for Computer Vision (CV) by utilizing its robustness in pattern recognition and reasoning (mirchandani2023). However, well-known LLM such as GPT and BERT can be used in forecasting time series data by fine-tuning it based on temporal data and also unifying their framework so that it can be used in diverse sequential task (classification, anomaly detection, forecasting, etc.) (Zhou et al., 2023). Time-LLM were developed in this year, a time series LLM model were developed to work with time-series data by reprogramming time series inputs into text prototype representations and guiding LLM’s reasoning by augmenting the input context with declarative prompts (Jin et al., 2024). It was proven competitive against many recently advanced time series forecasting method such as GPT4TS, DLinear, PatchTST, etc.
@@ 2.2 Data Mode Decomposition
Data mode decomposition is a technique used to decompose complex time series data into many data modes or Intrinsic Mode Function (IMF) with less complexity. Empirical Mode Decomposition (EMD) is one of data decomposition method that decomposes a signal into a set of intrinsic mode functions (IMFs) and a residual trend, capturing components of varying frequency ranges adaptively. EMD operates by iteratively sifting the data to identify and extract these IMFs, making it particularly effective for non-stationary and nonlinear time series data (Huang et al., 1998). However, EMD can be sensitive to noise and may suffer from mode mixing, where different modes are incorrectly combined.
To address those limitations, Variational Mode Decomposition (VMD) was introduced as a more robust data decomposition method. VMD decomposes a signal into modes by solving a variational problem that seeks to minimize the bandwidth of each mode while ensuring they are as distinct as possible from each other. It utilized Lagrange function and Alternate Direction Method of Multipliers (ADMM) (Dragomiretskiy & Zosso, 2014). Lagrange function uses $\lambda$ penalty term to enforce constrain for decomposed components. Lagrange is a quite complex function that can be simplified by implementing ADMM to split the function into several vital parts of function. ADMM works by iterating those functions until it reaches convergence (Boyd et al., 2011). This method is less prone to noise and mode mixing issues compared to EMD, as it uses a principled optimization approach to separate modes in the frequency domain. Dynamic Mode Decomposition (DMD) is an another data decomposition method that is capable of extracting important dynamic characteristics of the data. DMD decomposes the data into several data modes or IMFs by utilizing matrix analysis and singular value decomposition (Proctor et al., 2016).
@@ 3.1 VMD-Linear Framework
The structure of VMD-Linear shown in Figure 1 was inspired by VMD-BLSTM framework introduced by Zhang et al. (2021). The goal is to reduce data bandwidth in all IMF or data modes with the sum of all IMF equals to that of the original data. LTSF-Linear models will be assigned on each IMF for time-series forecasting. As the final step, the individual forecasts will be aggregated to produce the overall forecast result.
@@ 3.2 Variational Mode Decomposition
The symbol $\phi_{k}(t)$ is a non-decreasing function (can be considered a wave) with the condition that $\phi_{k}(t)\geq 0$, and $A_{k}(t)$ is the envelope of the wave $\phi_{k}(t)$. All IMF formed must satisfy Equation 1. VMD applies Hilbert transform to obtain the analytic frequency spectrum ($Au_{k}$) of each mode $u_{k}$. In the transformation process, each mode is multiplied by the center frequency ($e^{-jw_{k}t}$) to shift its frequency spectrum towards the baseband (Dragomiretskiy & Zosso, 2014).
@@ 3.3 Long-term Time Series Forecasting Linear
NLinear As shown in Figure 2, NLinear model is used to anticipate distribution shifts in the data by reducing the last input passed to the Linear layer and then adding back the reduced input before making the final prediction.
@@
DLinear DLinear enhances the simple linear model by utilizing moving average to split the data into Trend and Seasonal with each going through their seperate linear layer. This model should be used if there is a clear trend in the data. DLinear framework is shown in Figure 3.
@@ 4.1 Dataset
    Wind Turbine, The following data contains information about a certain wind turbine in Turkey that were captured using the SCADA system (Erisen2019). Wind Turbine related data such as wind speed, wind direction, and power generated were captured in 10-minute intervals with a total of 50,500 data points.
@@
    SEA Air Quality Dataset We also used various real-world Air Quality Index (AQI) datasets from many cities recorded from 27 January 2022 to 22 Mei 2023 (AQI2023). Those cities were Central Singapore, Banjarbaru, Batam, Jakarta, South Jakarta, Jambi, Malang, Medan, Samarinda, and Semarang.
@@ 4.2 Experimental Protocols
Dataset Linear DLinear NLinear LSTM BiLSTM RNN
@@
Dataset Linear DLinear NLinear LSTM BiLSTM RNN
@@ 5 Conclusion And Future Work
Further research should explore optimal forecasting models and try various preprocessing methods. LTSF-Linear has a limited capacity in handling long term time series forecasting but can be further improved by applying various preprocessing methods.